# Iola Walker: A Mobile Footfall Detection System for Music Composition

### v1

@@ Abstract
This project is the first of several experiments composing music that changes in response to biosignals. The system is dubbed "iola walker" in reference to a common polyrhythm, the hemiola. A listener goes for a walk, and the Iola Walker app detects their walking pace. Iola Walker picks up footfalls using a foot-mounted accelerometer, processing the signals in real time using a recurrent neural network [3] in an android app. The android app outputs a midi event for each footfall. The iola walker player, which might be a VST running in a DAW, plays the version of the next music passage with underlying polyrhythms closest to the listener’s walking pace. This paper documents the process of training the model to detect the footfalls in real time. The model is trained on accelerometer data from an Mbient Labs foot-mounted IMU [1] at 200 Hz, with the ground truth for footfalls annotated by pressing the volume up button on the android device when the foot hits the ground. To collect training data, I walked around my neighborhood clicking the volume up button each time my foot hit the ground. Several methods were tried for detecting footfalls in real time from sensor data, including ones based on digital signal processing techniques and traditional machine learning techniques.
I Android Data Acquisition Pipeline
@@
[…]
@@
II Baseline Models and the Journey to the Transition to Conv–LSTM
Before arriving at our final Conv–LSTM architecture, we evaluated two simpler approaches—a 1D convolutional neural network (CNN) and a logistic regression classifier—using the same sliding-window dataset described in Section 2. Although both baselines trained successfully, their performance rankings (logistic regression $<$ CNN $<$ Conv–LSTM) motivated our shift to the LSTM model.
II-A 1D Convolutional Neural Network
@@ II-B Logistic Regression
While extremely fast to train, logistic regression lacked the representational power to capture the temporal structure of footfalls, yielding the lowest F1 score.
II-C Comparative Results
@@
[…]
III Conv–LSTM Data Preparation and Script Configuration
@@ III-A Dataset Description
III-B Entry-Point Script Overview
@@
III-C Configuration Constants
At the top of the script we define all hyper-parameters and file paths in one place, enabling rapid experimentation:
@@ III-C Configuration Constants
III-D CSV Parsing and Label Assignment
@@ III-E Feature Engineering
III-F Sliding-Window Construction
@@
[…]
@@
[…]
@@
With data now loaded, labeled, scaled, and windowed, the pipeline proceeds to the training loop described in Section LABEL:sec:training_pipeline.
IV Conv–LSTM Model Definition
In this section we break down each component of the ConvLSTMGlobal architecture in detail, so that readers with minimal machine-learning background can follow.
IV-A Input Representation
@@
These are arranged as a tensor of shape $\bigl{(}B,\,W,\,F\bigr{)}$ for a batch of size $B$. We transpose this to $(B,\,F,\,W)$ before the 1D convolutions.
IV-B 1D Convolutional Layers
@@ IV-C Bidirectional LSTM
self.lstm = nn.LSTM(
@@
[…]
@@
IV-D Linear Heads for Prediction
@@
self.seq_head = nn.Linear(hid*2, 1) # per-timestep prediction
self.global_head = nn.Linear(hid*2, 1) # whole-window prediction
@@ V Training Pipeline
[…]
    • Compute two BCEWithLogitsLoss terms with analytically set pos_weight equal to #$\frac{\#\text{neg}}{\#\text{pos}}$ for sequence and global heads, then sum:
[…]
@@
[…]
@@
    • Trace the model to TorchScript on CPU using a zero tensor [4] of shape $(1,W,F)$ and save the serialized .pt file.
[…]
@@
VI Model Export and Metadata Packaging
@@
VII Evaluation: Tolerance Sweep Tester
@@
VIII Results Summary
VIII-A Validation Curve
@@
VIII-B Test-Set Tolerance Sweep
@@
Allowing a practical timing tolerance (±75 ms) yields a usable precision–recall trade-off.
IX Android Real-Time Inference Pipeline (Expanded)
@@
IX-A Model and Metadata Initialization
@@
IX-B Streaming Feature Buffer
@@ IX-C Model Inference
  • Flatten & normalize: Copies the deque into the pre-allocated buffer in $[1\times W\times 5]$ layout, then applies $x^{\prime}=\frac{x-\text{mean}}{\text{scale}}$ per channel.
@@
IX-D Decision Logic and Smoothing
@@
  • A sliding buffer of the last hitWindow boolean decisions contains at least requiredHits positives, implementing an $N$-of-$M$ majority filter to suppress isolated spikes.
@@
IX-E Runtime Parameter Tuning via Sliders
@@
These sliders live in the "Footfall Tuning" activity, updating the service’s state flows in real time so that you can walk, watch a live confidence graph, and immediately hear the impact of different settings.
IX-F Output Actions
Upon confirming a footfall, the service:
  • Emits an OSC message tagged /footfall with the precise timestamp for downstream audio or visual effects.
  • Plays a 50 ms "ding" via SoundPool (or ToneGenerator fallback).
  • Updates a Kotlin StateFlow counter, driving an on-screen display of total footfalls detected.
IX-G Performance Characteristics
On a mid-range Android device:
  • Inference time: $\sim$7 ms per window, which is 0.5 s long.
  • Latency: end-to-ding at 0.5 s (being improved upon).
  • Power draw: $\approx$3battery drain per hour during continuous walking.
The slider helps actually make the thing work; eventually it’ll be incorporated to work automatically.
X Conclusion and Future Work
Sometimes a working prototype is worth a thousand papers. I’d hope that the reader installs and tests the android app. There will need to be upgrades made in the future, in particular, writing the inference code in C++. Any audio processing code is going to require this. Here are specific upgrades that will be made in the near future.
  • Native C++ Inference Engine: Port the Python/TorchScript inference to C++ using the Android NDK. This eliminates JNI and Java-level overhead, further reducing latency and jitter.
  • rtneural Integration: Replace the heavy LSTM implementation with rtneural, a header-only C++ library for neural networks optimized for real-time audio. By exporting the trained Conv–LSTM weights into rtneural layers, the model can run entirely within the audio thread.
  • Low-Level Audio Callbacks: Integrate inference directly into the Oboe or OpenSL ES audio callback, so that footfall events can trigger sample-accurate MIDI or OSC messages without thread handoffs.
  • MIDI/OSC Output to Plugins: Use the Android USB MIDI API or built-in OSC client to deliver each detected footfall as a timestamped MIDI note or OSC message to a VST/AU plugin running on the device. This will allow artists to map footfalls to percussive samples, effects parameters, or synth envelopes with sample-level precision.
  • Adaptive Model Updates: Gather live performance data (timestamps, confidence values, false-positive logs) and periodically re-train or fine-tune the model offline. This continuous calibration can compensate for changes in walking surface, footwear, or sensor placement.
There were a few engineering challenges that needed to be addressed in this project. With online advice, training the footfall detection model in Python was substantially easier than it was even a year ago. The hard part was debugging in the physical world with real signals. Our AI codewriting companions still do not understand as well as we what it is like to operate in a physical body with physical objects.

### v4

@@ Abstract
This outing is part of a larger music technology research project. The objective is to find a method for materially enhancing music using hardware and software. There is a strong likelihood that there exists a new medium for experiencing music via a wearable device that ordinary listeners prefer over the current state of the art. If such a medium is discovered, it is a step towards altruistic, prosocial reform in the music industry. A new playback system infrastructure has a chance to soothe some of the societal problems tied to the larger entertainment industry ecosystem. Artifacts for this paper are available here: https://github.com/willbjames/iolawalker
I Introduction: How the new Music Composition and Listening System Works
The project is dubbed "iola walker" in reference to a common polyrhythm, the hemiola. To create an interactive listening experience for the wearer, musicians record songs with a walking listener in mind, imagining a range of walking paces while playing rhythmic parts. The same passages in a musical composition might be recorded multiple times with a variety of underlying pulses. In the associated interactive listening experience, a listener dons a foot-mounted inertial measurement unit and goes for a walk. The Iola Walker mobile app detects their footfalls in real time from the IMU over bluetooth.
Novel music composition gestures can be derived and implemented using this system. Perhaps a playback system could select the version of a recorded musical passage with underlying polyrhythms or "pulse" closest to the listener’s walking pace. Perhaps the crash at the end of a drum fill could be timed to line up precisely with a footstep. Perhaps the next passage of a song could be determined according to a detected change in walking speed that may have been induced by the previous section.
Information about predicted future footfalls is passed from the footfall detection process to the audio playback code running in a render process. This design allows for relevant signals, such as footfalls, to be incorporated into music with zero latency.
This paper documents the process of training a model to detect a walking listener’s footfalls from real-time signal data, as well as exporting that model to work with low latency on an android application. The model was trained on signals from an Mbient Labs foot-mounted IMU [1] collecting data at 200 Hertz, with the ground truth for footfalls annotated by pressing the volume up button on a BLU JL8 android device [2] when the listener’s foot hits the ground. To collect training data, I walked around my neighborhood while clicking the volume up button on an android mobile phone each time my foot hit the ground. I tried several models, and found that an LSTM attained the best performance.
II Related Works: The State of The Art
The signal from a person’s gait has been identified by other teams as a strong candidate for music technology research. But prior research on gait-modulated music ended up focusing on what the wearer was doing more than the music being played.
The Djogger research project was the first to explore this technology in order to help enhance exercise in 2006 [3]. Then several startups in the 2010s including Weav and Rock My Run, created a system for running workouts by changing the speed of songs played to match running pace [4, 5]. Most recently, in the 2020s, Medrhythms, in a partnership with Universal Music Group, has obtained FDA approval for its InTandem music therapy technology to be prescribed for people recovering from a stroke. [6, 7].
This project is different because the focus is not the wearer’s body motions. This project is decidedly not intended to develop a workout aid. It is a method for music composers to augment their art.
Our method to detect the footfall signal so that it is usable in real time is similar to prior LTSM methods documented as being highly accurate [8, 9, 10]. Our training and inference implementation has several unique features stemming from the intention to use it for music. The way to train a model to output a footfall control signal for an audio rendering process is documented in this work.
III Mobile Data Acquisition Pipeline
@@
  • Background thread that drains the queue, batches samples, and flushes to a $64\text{\,}\mathrm{kB}$ buffered CSV writer every $1\text{\,}\mathrm{s}$ or 1000 samples.
@@
IV The path to Conv–LSTM
Before arriving at our final Conv–LSTM architecture, we evaluated two simpler approaches, a 1D convolutional neural network and a logistic regression classifier, using the same sliding-window dataset described in Section 2. Although both baselines trained successfully, only the LSTM was accurate enough for real-time inference.
IV-A 1D Convolutional Neural Network
@@ IV-B Logistic Regression
While fast to train, logistic regression lacked the representational power necessary for time series data of footfalls, yielding the lowest F1 score.
IV-C Comparative Results
@@
$\text{F}_{1}(\text{logistic})<\text{F}_{1}(\text{CNN})<\text{F}_{1}(\text{Conv--LSTM})$
V Conv–LSTM Data Preparation and Script Configuration
@@ V-A Dataset Description
V-B Entry-Point Script Overview
@@
V-C Configuration Constants
At the top of the script are defined all hyper-parameters and file paths in one place, facilitating quick experimentation:
@@ V-C Configuration Constants
V-D CSV Parsing and Label Assignment
@@ V-E Feature Engineering
V-F Sliding-Window Construction
@@
After this loop, $\texttt{len(self.windows)}=\lfloor(N-W)/S\rfloor+1$ windows. The dataset’s __getitem__ returns:
@@
  • $\texttt{seq\_labels}\in\{0,1\}^{\text{WINDOW\_SIZE}}$,
  • $\texttt{glob\_label}\in\{0,1\}$.
@@
With data now loaded, labeled, scaled, and windowed, the pipeline proceeds to the training loop described in Section 5.
VI Conv–LSTM Model Definition
In this section we break down each component of the ConvLSTMGlobal architecture in detail.
VI-A Input Representation
@@
These are arranged as a tensor of shape $\bigl{(}B,\,W,\,F\bigr{)}$ for a batch of size $B$. We transpose this to (B,F,W)(B,\,F,\,W) before the 1D convolutions.
VI-B 1D Convolutional Layers
@@ VI-C Bidirectional LSTM
self.lstm = $n$.LSTM(
@@
  • $\texttt{hidden\_size}=hid$ controls the capacity of this memory (the larger, the more complex patterns it can learn, at the cost of more parameters).
@@
VI-D Linear Heads for Prediction
@@
self.seq_head = $n$.Linear(hid*2, 1) # per-timestep prediction
self.global_head = $n$.Linear(hid*2, 1) # whole-window prediction
@@ VII Training Pipeline
    • Forward through ConvLSTMGlobal: yields $\texttt{seq\_logits}\in\mathbb{R}^{B\times W}$ and $\texttt{glob\_logits}\in\mathbb{R}^{B}$.
    • Compute two BCEWithLogitsLoss terms with analytically set pos_weight equal to $\frac{\#\text{neg}}{\#\text{pos}}$ for sequence and global heads, then sum:
      $\mathcal{L}$ $=\frac{1}{B\,W}\sum\mathrm{BCE}(\texttt{seq\_logits},\texttt{seq\_targets})$
      $\quad+\frac{1}{B}\sum\mathrm{BCE}(\texttt{glob\_logits},\texttt{glob\_targets}).$
@@
  • Threshold sweep: After each epoch, disable gradients and run the model over the entire validation set to collect logits. For thresholds $\tau\in\{0.01,0.02,\dots,0.90\}$, compute F1 on binarized $\texttt{seq\_logits}>\tau$. Select the $\tau^{\star}$ maximizing F1.
@@
    • Trace the model to TorchScript on CPU using a zero tensor [paszke2019] of shape $(1,W,F)$ and save the serialized .pt file.
    • Write a JSON containing $\{W,S,\tau^{\star},\text{scaler\_mean},\text{scaler\_var}\}$ for mobile inference.
@@
VIII Model Export and Metadata Packaging
@@
IX Evaluation: Tolerance Sweep Tester
@@
X Results Summary
X-A Validation Curve
@@
X-B Test-Set Tolerance Sweep
@@
Allowing a practical timing tolerance (±75 ms) yields a usable precision–recall trade-off. We’ll need timing to be more precise eventually, but for a scenario where we’re choosing passages of music to play in the scale of a few seconds, it’s good enough for now.
XI Android Real-Time Inference Pipeline
@@
XI-A Model and Metadata Initialization
@@
XI-B Streaming Feature Buffer
@@ XI-C Model Inference
  • Flatten & normalize: Copies the deque into the pre-allocated buffer in $1\times W\times 5]$ layout, then applies $x^{\prime}=\frac{x-\text{mean}}{\text{scale}}$ per channel.
@@
XI-D Decision Logic and Smoothing
@@
  • A sliding buffer of the last hitWindow boolean decisions contains at least requiredHits positives, implementing an NN-of-$M$ majority filter to suppress isolated spikes.
@@
XI-E Runtime Parameter Tuning via Sliders
@@
These sliders live in the "Footfall Tuning" activity, updating the underlying parameters in real time.
XI-F Performance Characteristics
On a BLU JL8 Android device:
  • Inference time: $\sim$7 ms per window.
  • Latency: end-to-ding at 0.5 s.
The android app provides an example for a signal provider that could be consumed by an audio process. The information passed to the audio process would correspond to predetermined points in a song, such as beats. The main, strong footfall signal has been captured here with suitable latency for sending relevant composition information in a timely manner to the audio process.
XII Results and Discussion
As of the date of this release, the mobile app outputs a ding when a footfall is detected, in order to show what’s going on in real time. The footfall signal was successfully picked up with a low enough latency to be usable for music composition. With that said, existing musical protocols such as midi do not suffice to convey the information from the footfall detection control process to the audio playback rendering process.
The audio processing code needs to operate stably in real time, but the control code to pass the predicted footfall information can be more flexible in order to accomodate substantial inference methods like a bi-LSTM. The control code can then pass messages to the C++ audio code without needing to adhere to strict timing constraints. If this approach to predictive music composition turns out to be effective, it might be useful to implement a standard for passing information such as footfall signal between control and rendering process that iterates on Midi 2.0.
One difficulty of this system is that the methods for composing music are very interactive. Digital audio workstations and instruments are rich bidirectional sources of information. In order for this system to be effective, a representation of the internal walking state needs to be clear to the composer so they can design gestures in the music and record multiple parts based on how that signal changes. A better visualization or audiation method borrowed from audio engineering software could be implemented to interact with the footfall signal.
XIII Conclusions and Future Directions
One thing that became clear in doing this research is that music strongly skews towards being a one-way, unilateral form of communication. Specifically, it’s often going to be the most compelling move from an artistic standpoint to capture exactly what the person played in the moment. There are not many existing examples of highly effective, interactive musical compositions.
It’s a worthwhile challenge to create highly interactive music because it hasn’t been done effectively many times. But there’s definitely an increased risk of putting lipstick on a pig, in other words, spending time on a research direction that is unlikely to work.